# Mathematical Theorems on Turbulence

### v1

@@ 1. The Euler equations of ideal fluid motion
By Stokes theorem, this line integral is equivalent to the flux of the curl of the velocity vector field $u$, the vorticity $\omega$, through any comoving bounding surface. Letting the loop become infinitesimal about any given point $x\in M$, with arbitrary orientation, we recover the Lie advection equation for the vorticity. In two dimensions one can identify the vorticity with a scalar field $\omega=\nabla^{\perp}\cdot u$ where $\nabla^{\perp}=(-\partial_{2},\partial_{1})$ and in three dimensions with a vector field $\omega=\nabla\times u$. These are transported by $u$:
@@
$(\partial_{t}+u\cdot\nabla)\nabla u+(\nabla u)^{2}$ $=-\nabla^{2}p$ (1.10)
$-\Delta p$ $={\rm tr}(\nabla u)^{2}$ (1.11)
$\partial_{n}p|_{\partial M}$ $=u\cdot\nabla\hat{n}\cdot u|_{\partial M}$ (1.12)
@@
In fact, a flurry of recent progress indicates that even infinitely smooth initial conditions become singular, at least in the case where there is a solid boundary [35]. See discussion in [66]. Thus, turbulence – which we will argue lies outside the class of functions which result in a local theory – cannot be avoided with certainty, even if one starts from the non-turbulent regime. Of arguably more relevance is the fact that turbulence in real world scenarios should not be modeled as arising from regular initial conditions – rather it often results from initial conditions which are already much too rough for a predictive Cauchy problem to exist. We shall discuss this point in the subsequent section.
@@ 2. Navier-Stokes and the inviscid limit
where $\nu>0$ represents the kinematic viscosity of the medium, and $f$ an external force. Dry water corresponds to a material with $\nu=0$ exactly, but no such exists. However, it turns out the many flows that we encounter in Nature are almost dry! To understand this point, note that the force of this new viscous effect depends on the state of motion. Indeed, if – for example – we non-dimensionalize the system using a characteristic velocity $\mathsf{U}=\|u_{0}\|_{L^{2}}$, force $\mathsf{F}:=\|f\|_{L^{2}}$, length $\mathsf{L}={\rm diam}(M)$ and time $\mathsf{T}=\mathsf{L}/\mathsf{U}$, then the resulting non-dimensional equations now feature the Reynolds number $\mathsf{Re}=\mathsf{U}\mathsf{L}/\nu$ and (modified) Grashof number $\mathsf{Gr}:=\mathsf{F}/(\mathsf{U}^{2}/\mathsf{L})$Grashof number conventionally means $\widetilde{\mathsf{Gr}}:=\mathsf{F}L^{3}/\nu^{2}$, which mixes small scale with large scale quantities [43]. It is easy to see that $\widetilde{\mathsf{Gr}}=\mathsf{Gr}\mathsf{Re}^{2}$. The reason behind this definition is that it involves purely large scale quantities, and (assuming Taylors hypothesis $\langle{\varepsilon}\rangle=\mathsf{U}^{3}/L$) it equates to $\widetilde{\mathsf{Gr}}=\mathsf{F}\mathsf{U}/\langle{\varepsilon}\rangle$. Thus, we see that $\mathsf{Gr}$ measures the ”inefficiency” of the forcing function, becoming very large when $\langle{\varepsilon}\rangle\ll\mathsf{F}\mathsf{U}$, so that force injects energy inefficiently compared with a naive dimensional estimate. See [98, 146]. , two dimensionless parameters of the system
@@
$S_{p}^{u^{\nu}}(\ell):=\fint_{0}^{T}\fint_{M}\fint_{\mathbb{S}^{d-1}}|u^{\nu}(x+\ell\hat{z},t)-u^{\nu}(x,t)|^{p}\ {\rm d}\sigma(\hat{z}){\rm d}x{\rm d}t$ (2.7)
@@
Figure 3. Measurements of second order structure functions and their exponents from numerical simulation [65].
@@ 3. Onsager’s ideal turbulence
The result of the last section is that, under an observed uniform regularity assumption, the high Reynolds number limit of a Navier-Stokes solution is a locally dissipative weak solution of the Euler equation. Anomalous dissipation is the phenomenon that this local dissipation measure is non-trivial. Moreover, we see that it can be computed in two ways. First, it can be computed as an inviscid limit of the viscous local dissipation measure 2.18. But, evidently, it can also be computed on the limiting weak solution itself (2.21). As such, we now seek another way to compute the dissipation on the inviscid weak solution $u$ which illuminates its ultimate cause. Recall from our earlier considerations around (2.16) that we expect having exactly $\nicefrac{{1}}{{3}}$ of a derivative in $L^{3}_{t,x}$ to play a special role. The person to envisage irregular weak Euler solutions as describing turbulent dynamics, and to realize this precise regularity threshold was the Nobel Laureate Lars Onsager, see left panel of Figure 5. He wrote a now famous statement to this effect in his 1949 paper on Statistical Hydrodynamics [160]. There he stated that, in order for anomalous dissipation to take place, the fluid cannot satisfy a Hölderian condition of the form
@@
Figure 5. Lars Onsager and the anomaly [161, 99, Folder 11:129, p.14]
@@
Thus, the pressure $p\in L^{\frac{3}{2}}_{x,t}$ by Lemma 3.2. We now introduce mollification of the weak solution $\overline{u}_{\ell}$, on which we may perform manipulations. The mollification procedure was described already in the proof of Lemma 2.1, see Figure 6 for a visualization. Specifically, we choose $\overline{u}_{\ell}$ as our test function in weak formulation of the Euler/Navier-Stokes equations (2.13)–(2.14).
@@
$4{\varepsilon}_{\ell}[u]$ $=-\text{div}\overline{(|u|^{2}u)}_{\ell}+u\cdot\nabla\overline{(|u|^{2})}_{\ell}+2u\cdot\text{div}\overline{(u\otimes u)}_{\ell}-2u\otimes u:\nabla\overline{u}_{\ell}$
@@
Question 7.8 is particularly important because these spaces, in particular $u\in L^{3}_{t}B_{3,\infty}^{1/3}$, are natural candidates for the regularity of actual turbulent flows. We will discuss this further point in §4. For the moment, I remark that there is only one “critical space" (namely a space where the flux ${\varepsilon}_{\ell}[u]$ is bounded but does not vanish for general vector fields in that class) for which the answer to Question 7.8 is known. This critical space is $u\in L^{\infty}_{t,x}\cap L^{1}_{t}BV$, in which De Rosa and Inversi showed weak solutions conserve energy [57] (see also [118]). This result is remarkable since $L^{\infty}_{t,x}\cap L^{1}_{t}BV_{x}\subset L^{3}_{t}B_{3,\infty}^{1/3}$ sharply. Indeed, if $\|u(\cdot+r)-u(\cdot)\|_{L^{1}}\lesssim|r|$ and $\|u\|_{L^{\infty}}<\infty$, then by interpolation $\|u(\cdot+r)-u(\cdot)\|_{L^{p}}\lesssim|r|^{1/p}$ for all $p\geq 1$. Moreover for compressible models such as Burgers, shocks live in $u\in L^{\infty}_{t,x}\cap L^{1}_{t}BV_{x}$ and are dissipative. As for Question 7.8, I would guess the answer is no, although recent work [57] as well as [124] brings into question some of the intuition behind this guess. This will be discussed around Conjecture 4.5.
@@
$E^{\ell,\nu}$ $:=\frac{|{u^{\nu}-\overline{u^{\nu}}_{\ell}}|^{2}}{2}$ (3.17)
$Q^{\ell,\nu}$ $:=\left(\frac{|{u^{\nu}-\overline{u^{\nu}}_{\ell}}|^{2}}{2}+(p^{\nu}-\overline{p^{\nu}}_{\ell})\right)(u^{\nu}-\overline{u^{\nu}}_{\ell})$ (3.18)
$R^{\ell,\nu}$ $:=\overline{u^{\nu}}_{\ell}\otimes\overline{u^{\nu}}_{\ell}-\overline{(u^{\nu}\otimes u^{\nu})}_{\ell}$ (3.19)
@@
These are not difficult to show, see [83] for a clear discussion. Let us prove the estimates for $p<\infty$ as $p=\infty$ is simpler. Since $\int_{\mathbb{R}^{d}}G_{\ell}(z){\rm d}z=1$, we have $\overline{f}_{\ell}(x)-f(x)=\int_{\mathbb{R}^{d}}\left(f(x-z)-f(x)\right)G_{\ell}(z)\,{\rm d}z.$ By the Jensen’s inequality, we have
@@
where we used that ${\rm supp}G_{\ell}\subset B_{\ell}(0)$, as well as the following estimate
@@
Finally, one implication of this result, pointed out by Isett, is that the quality of energy dissipation for Euler flows is stable under perturbation, in the class $u\in L^{\infty}_{t}B^{1/3}_{3,\infty}$. Namely, in this class $\|\dot{e}\|_{L^{\infty}}\lesssim\|u\|_{L^{\infty}_{t}B^{1/3}_{3,\infty}}^{3}$ and $\dot{e}(t)$ varies continuously in $L^{\infty}$ as $u$ varies in the $L^{\infty}_{t}B^{\sigma}_{3,\infty}$ topology, namely $\|\dot{e}_{1}-\dot{e}_{2}\|_{L^{\infty}}\lesssim\|u_{1}-u_{2}\|_{L^{\infty}_{t}B^{1/3}_{3,\infty}}$. This means that anomalous dissipation $\dot{e}(t)\leq-{\varepsilon}<0$ is an open condition in the space of weak Euler solutions in the critical class endowed with the strong topology. By contrast, Isett conjectured that for weak solutions with regularity below $\nicefrac{{1}}{{3}}$ of a derivative, or with time integrability below $q=\infty$, the energy profile should generically fail to be of bounded variation on every time interval [120, 123]. In particular, a dissipative solution could be perturbed in this class to produce a nearby solution whose energy profile fails to be monotonic on every time interval. A version of this conjecture was proved by De Rosa and Tione [58] – namely the set of weak Euler solutions with energy profile not of bounded variation in any open interval is Baire generic in an appropriate complete metric space which is a strict subset of all $\alpha$–Hölder Euler solutions with $\alpha<1/3$.
@@ 4. Kolmogorov’s 1941 theory
We do not prove Proposition 4.2 here, as it unfortunately involves some cumbersome manipulations very similar to those which give rise to (3.3). Instead, we use Prop. 3.1 to establish a stronger result – the Kolmogorov $\nicefrac{{4}}{{3}}$ law in a finite range of scales
@@
$\qquad\quad\quad-\tfrac{1}{2}u^{\nu}\cdot(\overline{f}_{\ell}-f)-\tfrac{1}{2}(\overline{u}_{\ell}^{\nu}-u^{\nu})\cdot{f}$ (4.10)
in the sense of distributions. Since $u^{\nu}$ is compact in $L^{3}_{t,x}$, by the Frèchet–Kolmogorov compactness criterion, the quantity $\phi_{u}(\ell):=\sup_{|y|\leq\ell}\sup_{\nu>0}\|u^{\nu}(\cdot+y,\cdot)-u^{\nu}(\cdot,\cdot)\|_{L^{3}_{t,x}}$ defines a modulus of continuity, i.e. $\lim_{\ell\rightarrow 0}\phi_{u}(\ell)=0$. Moreover, a direct consequence of the Calderón–Zygmund estimate Lemma 3.2 is
@@
For example, in [69] it was proved that if $u^{\nu}\in L^{3}(0,T;B_{3,\infty}^{\sigma_{3}}(\mathbb{T}^{d}))$ uniformly, then $(\varepsilon^{\nu}[u^{\nu}],\psi)\leq\nu^{\frac{3\sigma_{3}-1}{\sigma_{3}+1}}$ (Corollary 3.12). When $\sigma_{3}>1/3$, this gives a bound on the rate $\alpha$. Our theorem shows a long scale range where the $\nicefrac{{4}}{{3}}$ law $S_{3,\rm mix}^{u^{\nu}}(\ell)\approx-\frac{4}{d}\varepsilon^{\nu}[u^{\nu}]\ell$ (similarly the $\nicefrac{{4}}{{5}}$ law $S_{3,\|}^{u^{\nu}}(\ell)\approx-\frac{12}{d(d+2)}\varepsilon^{\nu}[u^{\nu}]\ell$) is observed despite formally there not existing “anomalous dissipation" in the limit $\nu\to 0$. This is our interpretation of recent numerical data [124], from which Figure 10 is drawn. The relevant inertial range scales are
@@
The intuition behind this conjecture is the following. Assume anomalous dissipation occurs $\langle\varepsilon\rangle>0$. Then, Theorem 4.3 says that $\langle S_{3,\|}^{u^{\nu}}(\ell)\rangle/\ell$ is finite over a finite scale range, and non-vanishing as $\ell\to 0$ (provided it is taken in the "inertial range"). This range extends to zero length as $\nu\to 0$. Now, for any vector field in $L_{t}^{3}B_{3,\infty}^{s}$ with $s\geq 1/3$, the $\langle S_{3,\rm mix}^{u^{\nu}}(\ell)\rangle/\ell$ is bounded uniform in $\ell$ but if $s>1/3$ then the ratio clearly vanishes as $\ell\to 0$. Therefore we should have $s\leq 1/3$. But "generic" vector fields with $s<1/3$ should have $|\langle S_{3,\|}^{u}(\ell)\rangle/\ell\to\infty|$ as $\ell\to 0$, since $\langle S_{3,\|}^{u}(\ell)\rangle/\ell\sim(\delta_{\ell}u)^{3}/\ell\sim\ell^{3s-1}$. Turbulence should place the fluid velocity in such a generic category, why not? It thus seems that $L_{t}^{3}B_{3,\infty}^{1/3}$ is the only reasonable space, at least in the Besov scale. See also [65].
@@
Aside from the clearly shaky reasoning behind Conjecture 4.5, the recent numerical results obtained in [124] call into question some components of the intuition, namely the answer to Question 7.8 being negative. In particular, [124] gives evidence that the dissipation slowly vanishing in the inviscid limit. In this case, the information in the $\nicefrac{{4}}{{5}}$ law degenerates, and does not shed any particular light on the regularity of the limiting object. In fact, those numerics indicate the $L^{3}$ Besov regularity tends towards a value larger that $\nicefrac{{1}}{{3}}$ asymptotically which, via Corollary 3.12, is consistent with the perceived rate of decay of energy dissipation. See Figure 7. This may be taken as an (admittedly very weak) point in favor of the answer to Question 7.8 being positive, since why would the turbulent solution land on any particular $\sigma_{3}>1/3$?

### v2

@@ 1. The Euler equations of ideal fluid motion
By Stokes’ theorem, this line integral is equivalent to the flux of the curl of the velocity vector field $u$, the vorticity $\omega$, through any comoving bounding surface. Letting the loop become infinitesimal about any given point $x\in M$, with arbitrary orientation, we recover the Lie advection equation for the vorticity. In two dimensions one can identify the vorticity with a scalar field $\omega=\nabla^{\perp}\cdot u$ where $\nabla^{\perp}=(-\partial_{2},\partial_{1})$ and in three dimensions with a vector field $\omega=\nabla\times u$. These are transported by $u$:
@@
$(\partial_{t}+u\cdot\nabla)\nabla u+(\nabla u)^{2}$ $=-\nabla^{2}p,$ (1.10)
$-\Delta p$ $={\rm tr}(\nabla u)^{2},$ (1.11)
$\partial_{n}p|_{\partial M}$ $=u\cdot\nabla\hat{n}\cdot u|_{\partial M},$ (1.12)
@@
In fact, a flurry of recent progress indicates that even infinitely smooth initial conditions become singular, at least in the case where there is a solid boundary [35], see discussion in [66]. Thus, turbulence – which we will argue lies outside the class of functions which result in a local theory – cannot be avoided with certainty, even if one starts from the non-turbulent regime. Of arguably more relevance is the fact that turbulence in real world scenarios should not be modeled as arising from regular initial conditions – rather it often results from initial conditions which are already much too rough for a predictive Cauchy problem to exist. We shall discuss this point in the subsequent section.
@@ 2. Navier-Stokes and the inviscid limit
where $\nu>0$ represents the kinematic viscosity of the medium, and $f$ an external force. Dry water corresponds to a material with $\nu=0$ exactly, but no such exists. However, it turns out the many flows that we encounter in Nature are almost dry! To understand this point, note that the force of this new viscous effect depends on the state of motion. Indeed, if – for example – we non-dimensionalize the system using a characteristic velocity $\mathsf{U}=\|u_{0}\|_{L^{2}}$, force $\mathsf{F}:=\|f\|_{L^{2}}$, length $\mathsf{L}={\rm diam}(M)$ and time $\mathsf{T}=\mathsf{L}/\mathsf{U}$, then the resulting non-dimensional equations now feature the Reynolds number $\mathsf{Re}=\mathsf{U}\mathsf{L}/\nu$ and (modified) Grashof number $\mathsf{Gr}:=\mathsf{F}/(\mathsf{U}^{2}/\mathsf{L})$Grashof number conventionally means $\widetilde{\mathsf{Gr}}:=\mathsf{F}L^{3}/\nu^{2}$, which mixes small scale with large scale quantities [43]. It is easy to see that $\widetilde{\mathsf{Gr}}=\mathsf{Gr}\mathsf{Re}^{2}$. The reason behind this definition is that it involves purely large scale quantities, and (assuming Taylor’s hypothesis $\langle{\varepsilon}\rangle=\mathsf{U}^{3}/L$), where $\langle{\varepsilon}\rangle$ is the mean dissipation (2.18), it equates to $\widetilde{\mathsf{Gr}}=\mathsf{F}\mathsf{U}/\langle{\varepsilon}\rangle$. Thus, we see that $\mathsf{Gr}$ measures the ”inefficiency” of the forcing function, becoming very large when $\langle{\varepsilon}\rangle\ll\mathsf{F}\mathsf{U}$, so that force injects energy inefficiently compared with a naive dimensional estimate. See [98, 147]. , two dimensionless parameters of the system
@@
$S_{p}^{u^{\nu}}(\ell):=\fint_{0}^{T}\fint_{M}\fint_{\mathbb{S}^{d-1}}|u^{\nu}(x+\ell\hat{z},t)-u^{\nu}(x,t)|^{p}\ {\rm d}\sigma(\hat{z}){\rm d}x{\rm d}t,$ (2.7)
@@
Figure 3. Measurements of second order structure functions and their exponents from numerical simulation [65]. Right panel is best fit measured exponent $\zeta_{2}^{\nu}:=\log S_{2}^{u^{\nu}}(\ell)/\log\ell$ for $\ell$ in the inertial range.
@@ 3. Onsager’s ideal turbulence
The result of the last section is that, under an observed uniform regularity assumption, the high Reynolds number limit of a Navier-Stokes solution is a locally dissipative weak solution of the Euler equation. Anomalous dissipation is the phenomenon that this local dissipation measure is non-trivial. Moreover, we see that it can be computed in two ways. First, it can be computed as an inviscid limit of the viscous local dissipation measure (2.18). But, evidently, it can also be computed on the limiting weak solution itself, see (2.21). As such, we now seek another way to compute the dissipation on the inviscid weak solution $u$ which illuminates its ultimate cause. Recall from our earlier considerations around (2.16) that we expect having exactly $\nicefrac{{1}}{{3}}$ of a derivative in $L^{3}_{t,x}$ to play a special role. The person to envisage irregular weak Euler solutions as describing turbulent dynamics, and to realize this precise regularity threshold was the Nobel Laureate Lars Onsager, see left panel of Figure 5. He wrote a now famous statement to this effect in his 1949 paper on Statistical Hydrodynamics [161]. There he stated that, in order for anomalous dissipation to take place, the fluid cannot satisfy a Hölderian condition of the form
@@
Figure 5. Lars Onsager and the anomaly [162, 99, Folder 11:129, p.14]
@@
Thus, the pressure enjoys the regularity $p\in L^{\frac{3}{2}}_{x,t}$ by Lemma 3.2. We now introduce mollification of the weak solution $\overline{u}_{\ell}$, on which we may perform manipulations. The mollification procedure was described already in the proof of Lemma 2.1, see Figure 6 for a visualization. Specifically, we choose $\overline{u}_{\ell}$ as our test function in weak formulation of the Euler/Navier-Stokes equations (2.13)–(2.14).
@@
$4{\varepsilon}_{\ell}[u]$ $=-\text{div}\overline{(|u|^{2}u)}_{\ell}+u\cdot\nabla\overline{(|u|^{2})}_{\ell}+2u\cdot\text{div}\overline{(u\otimes u)}_{\ell}-2u\otimes u:\nabla\overline{u}_{\ell}.$
@@
Question 3.7 is particularly important because these spaces, in particular $u\in L^{3}_{t}B_{3,\infty}^{1/3}$, are natural candidates for the regularity of actual turbulent flows. We will discuss this further point in §4. For the moment, I remark that there is only one “critical space" (namely a space where the flux ${\varepsilon}_{\ell}[u]$ is bounded but does not vanish for general vector fields in that class) for which the answer to Question 3.7 is known. This critical space is $u\in L^{\infty}_{t,x}\cap L^{1}_{t}BV$, in which De Rosa and Inversi showed weak solutions conserve energy [57] (see also [118]). This result is remarkable since $L^{\infty}_{t,x}\cap L^{1}_{t}BV_{x}\subset L^{3}_{t}B_{3,\infty}^{1/3}$ sharply. Indeed, if $\|u(\cdot+r)-u(\cdot)\|_{L^{1}}\lesssim|r|$ and $\|u\|_{L^{\infty}}<\infty$, then by interpolation $\|u(\cdot+r)-u(\cdot)\|_{L^{p}}\lesssim|r|^{1/p}$ for all $p\geq 1$. Moreover for compressible models such as Burgers, shocks live in $u\in L^{\infty}_{t,x}\cap L^{1}_{t}BV_{x}$ and are dissipative. As for Question 3.7, I would guess the answer is no, although recent work [57] as well as [124] brings into question some of the intuition behind this guess. This will be discussed around Conjecture 4.5.
@@
$E^{\ell,\nu}$ $:=\frac{|{u^{\nu}-\overline{u^{\nu}}_{\ell}}|^{2}}{2},$ (3.17)
$Q^{\ell,\nu}$ $:=\left(\frac{|{u^{\nu}-\overline{u^{\nu}}_{\ell}}|^{2}}{2}+(p^{\nu}-\overline{p^{\nu}}_{\ell})\right)(u^{\nu}-\overline{u^{\nu}}_{\ell}),$ (3.18)
$R^{\ell,\nu}$ $:=\overline{u^{\nu}}_{\ell}\otimes\overline{u^{\nu}}_{\ell}-\overline{(u^{\nu}\otimes u^{\nu})}_{\ell},$ (3.19)
@@
These are not difficult to show, see [83] for a clear discussion. Let us prove the estimates for $p<\infty$ as $p=\infty$ is simpler. Since $\int_{\mathbb{R}^{d}}G_{\ell}(z){\rm d}z=1$, we have $\overline{f}_{\ell}(x)-f(x)=\int_{\mathbb{R}^{d}}\left(f(x-z)-f(x)\right)G_{\ell}(z)\,{\rm d}z.$ By Jensen’s inequality, we have
@@
where we used that ${\rm supp}(G_{\ell})\subset B_{\ell}(0)$, as well as the following estimate
@@
Finally, one implication of this result, pointed out by Isett, is that the quality of energy dissipation for Euler flows is stable under perturbation, in the class $u\in L^{\infty}_{t}B^{1/3}_{3,\infty}$. Namely, in this class $\|e^{\prime}\|_{L^{\infty}}\lesssim\|u\|_{L^{\infty}_{t}B^{1/3}_{3,\infty}}^{3}$ and $e^{\prime}(t)$ varies continuously in $L^{\infty}$ as $u$ varies in the $L^{\infty}_{t}B^{\sigma}_{3,\infty}$ topology, namely $\|e^{\prime}_{1}-e^{\prime}_{2}\|_{L^{\infty}}\lesssim\|u_{1}-u_{2}\|_{L^{\infty}_{t}B^{1/3}_{3,\infty}}$. This means that anomalous dissipation $e^{\prime}(t)\leq-{\varepsilon}<0$ is an open condition in the space of weak Euler solutions in the critical class endowed with the strong topology. By contrast, Isett conjectured that for weak solutions with regularity below $\nicefrac{{1}}{{3}}$ of a derivative, or with time integrability below $q=\infty$, the energy profile should generically fail to be of bounded variation on every time interval [120, 123]. In particular, a dissipative solution could be perturbed in this class to produce a nearby solution whose energy profile fails to be monotonic on every time interval. A version of this conjecture was proved by De Rosa and Tione [58] – namely the set of weak Euler solutions with energy profile not of bounded variation in any open interval is Baire generic in an appropriate complete metric space which is a strict subset of all $\alpha$–Hölder Euler solutions with $\alpha<1/3$.
@@ 4. Kolmogorov’s 1941 theory
We do not prove Proposition 4.2 here, as it unfortunately involves some cumbersome manipulations very similar to those which give rise to (3.3). Instead, we use Proposition 3.1 to establish a stronger result – the Kolmogorov $\nicefrac{{4}}{{3}}$ law in a finite range of scales
@@
$\qquad\quad\quad-\tfrac{1}{2}u^{\nu}\cdot(\overline{f}_{\ell}-f)-\tfrac{1}{2}(\overline{u}_{\ell}^{\nu}-u^{\nu})\cdot{f},$ (4.10)
in the sense of distributions. Since $u^{\nu}$ is compact in $L^{3}_{t,x}$, by the Frèchet–Kolmogorov compactness criterion, the quantity $\phi_{u}(\ell):=\sup_{|y|\leq\ell}\sup_{\nu>0}\|u^{\nu}(\cdot+y,\cdot)-u^{\nu}(\cdot,\cdot)\|_{L^{3}_{t,x}}$ defines a modulus of continuity, i.e. $\lim_{\ell\rightarrow 0}\phi_{u}(\ell)=0$. Moreover, a direct consequence of the Calderón–Zygmund estimate in Lemma 3.2 is
@@
For example, in [69] it was proved that if $u^{\nu}\in L^{3}(0,T;B_{3,\infty}^{\sigma_{3}}(\mathbb{T}^{d}))$ uniformly, then $(\varepsilon^{\nu}[u^{\nu}],\psi)\leq\nu^{\frac{3\sigma_{3}-1}{\sigma_{3}+1}}$ (Corollary 3.12). When $\sigma_{3}>1/3$, this gives a bound on the rate $\alpha$. Our theorem shows a long scale range where the $\nicefrac{{4}}{{3}}$ law $S_{3,\rm mix}^{u^{\nu}}(\ell)\approx-\frac{4}{d}\varepsilon^{\nu}[u^{\nu}]\ell$ (similarly the $\nicefrac{{4}}{{5}}$ law $S_{3,\|}^{u^{\nu}}(\ell)\approx-\frac{12}{d(d+2)}\varepsilon^{\nu}[u^{\nu}]\ell$) is observed despite formally there is not existing “anomalous dissipation" in the limit $\nu\to 0$. This is our interpretation of recent numerical data [124], from which Figure 10 is drawn. The relevant inertial range scales are
@@
The intuition behind this conjecture is the following. Assume anomalous dissipation occurs $\langle\varepsilon\rangle>0$. Then, Theorem 4.3 says that $\langle S_{3,\|}^{u^{\nu}}(\ell)\rangle/\ell$ is finite over a finite (inertial) scale range. This range extends to zero length as $\nu\to 0$, and $\langle S_{3,\|}^{u^{\nu}}(\ell)\rangle/\ell$ is non-vanishing as $\ell\to 0$ . Now, for any vector field in $L_{t}^{3}B_{3,\infty}^{s}$ with $s\geq 1/3$, the $\langle S_{3,\rm mix}^{u^{\nu}}(\ell)\rangle/\ell$ is bounded uniform in $\ell$ but if $s>1/3$ then the ratio clearly vanishes as $\ell\to 0$. Therefore we should have $s\leq 1/3$. But "generic" vector fields with $s<1/3$ should have $|\langle S_{3,\|}^{u}(\ell)\rangle/\ell\to\infty|$ as $\ell\to 0$, since $\langle S_{3,\|}^{u}(\ell)\rangle/\ell\sim(\delta_{\ell}u)^{3}/\ell\sim\ell^{3s-1}$. Turbulence should place the fluid velocity in such a generic category, why not? It thus seems that $L_{t}^{3}B_{3,\infty}^{1/3}$ is the only reasonable space, at least in the Besov scale. See also [65].
@@
Aside from the clearly shaky reasoning behind Conjecture 4.5, the recent numerical results obtained in [124] call into question some components of the intuition, namely the answer to Question 3.7 being negative. In particular, [124] gives evidence that the dissipation slowly vanishing in the inviscid limit. In this case, the information in the $\nicefrac{{4}}{{5}}$ law degenerates, and does not shed any particular light on the regularity of the limiting object. In fact, those numerics indicate the $L^{3}$ Besov regularity tends towards a value larger that $\nicefrac{{1}}{{3}}$ asymptotically which, via Corollary 3.12, is consistent with the perceived rate of decay of energy dissipation. See Figure 7. This may be taken as an (admittedly very weak) point in favor of the answer to Question 3.7 being positive, since why would the turbulent solution land on any particular $\sigma_{3}>1/3$?